Evaluate integration integration X(y-1) dA where R is the region bounded by y=1-x^2 and y=x^2-3
1
Expert's answer
2021-07-20T08:10:28-0400

"\iint x(y-1)dR," where "R=\{y=1-x^2, y= x^2-3\}"

First find intersepts points of "y = 1-x^2" and "y = x^2-3":

"1-x^2 = x^2-3\\
2x^2 = 4\\
x^2 = 2\\
x = \pm \sqrt2"

So "\iint x(y-1)dR = \int_{-\sqrt2}^{\sqrt 2}dx \int_{x^2-3}^{1-x^2}x(y-1)dy ="

"= \int_{-\sqrt2}^{\sqrt 2}xdx(\cfrac{y^2}{2}-y|_{x^2-3}^{1-x^2}) = \\
=\int_{-\sqrt2}^{\sqrt 2}x(\cfrac{(1-x^2)^2}{2}-(1-x^2)-(\cfrac{(x^2-3)^2}{2}-(x^2-3)))dx=\\
=\int_{-\sqrt2}^{\sqrt 2}x(\cfrac{4x^2-8}{2}-4+2x^2)dx = \\
= \int_{-\sqrt2}^{\sqrt 2}x(4x^2-8)dx = \int_{-\sqrt2}^{\sqrt 2}(4x^3-8x)dx = x^4 - 4x^2 |_{-\sqrt 2}^{\sqrt 2} =\\
=4 - 4*2 -(4-4*2) = 0"